# Improved Treatment of 1-4 interactions in Force Fields for Molecular Dynamics Simulations

### v2

@@ 1 Introduction
The more common approach, employed by force fields such as OPLS 1, 2, AMBER 3, 4, CHARMM 5, 6, the Open Force Field class of FFs 7, 8, and AMOEBA9, 10, is to use a single scaling factor to modulate 1-4 non-bonded interactions. While this method is computationally simple and easy to implement, it is suboptimal for a diverse set of molecular interactions for several reasons. A single scaling parameter cannot capture the nuanced physical phenomena governing 1-4 interactions across different chemical environments, leading to reduced accuracy in force field predictions. While GROMOS 11, 12 incorporates unique 1-4 parameters in specific cases, it defaults to scaled interactions for most systems. While this tailored approach has the potential to improve accuracy for certain molecules or functional groups, it does not fully address the underlying limitations of the standard Lennard-Jones and Coulomb potentials. These functional forms inherently lack the correct physics to represent interactions at short distances characteristic of 1-4 pairs. Although torsional barriers can be improved by fine-tuning 1-4 parameters, this does not guarantee accurate forces or a well-behaved potential energy surface (PES), which are critical for reliable simulations.
@@
To address the inherent limitations of traditional methods for handling 1-4 interactions, we propose two possible pathways for improvement. The first involves treating 1-4 interactions entirely through non-bonded potentials with physically accurate functional forms. This would necessitate the development or adoption of advanced force fields where all non-bonded interactions, including those at short distances, are damped using physically motivated corrections, such as density overlap approaches 13, 14, 15. Force fields like HIPPO 16 and CMM 17, which incorporate density overlap to account for charge penetration and related effects, are promising candidates for this strategy. Such an approach could significantly reduce the reliance on torsional terms by inherently capturing the correct balance of interactions, including electrostatics, van der Waals forces, polarization, and charge transfer. However, while this approach holds promise, achieving the correct balance among these interaction components remains a considerable challenge. Subtle inaccuracies in modeling electrostatics, van der Waals interactions, or polarization effects could introduce errors that may still need to be compensated for by torsional terms. Additionally, the importance of polarization effects, particularly for 1-2 and 1-3 interactions, cannot be overlooked, as they can contribute significantly to the overall polarization energy at the 1-4 level 18. Therefore, even with an advanced non-bonded potential, some torsional corrections may still be necessary to address residual discrepancies and ensure an accurate representation of the energy landscape.
@@
The alternative approach involves treating the 1-4 interactions entirely as bonded terms. This bonded-only methodology not only circumvents the challenges associated with charge penetration effects discussed earlier but also decouples the non-bonded and torsional parameterization processes. This separation allows torsional terms to be directly optimized against quantum mechanical (QM) reference data without interference from non-bonded interactions. Such an approach, however, necessitates handling the coupling between 1-4 interactions and 1-2 or 1-3 interactions—–effects that are traditionally, albeit imperfectly, addressed by non-bonded interactions in standard force fields. Fortunately, extensive research on coupled interaction terms, such as torsion-bond and torsion-angle coupling, provides a foundation for addressing these challenges. These terms were developed in the 1990s and implemented in force fields such as MM3 19, 20 and MMFF94 21, 22, CFF93 23, 24. However, to our knowledge, all existing implementations still rely on non-bonded 1-4 interactions in addition to the coupling terms. Moreover, despite their potential, these coupling terms have so far not been widely adopted in common force fields. We believe the limited adoption of these coupling terms can be attributed to two primary challenges. First, there has been a lack of tools capable of systematically parameterizing this plethora of coupling terms. Second, generating the necessary QM reference data to calibrate these parameters accurately can also be a significant hurdle. Addressing these gaps requires an automated framework capable of deriving and optimizing these terms efficiently, as well as creating the necessary QM datasets.
@@
In this work, we revisit the earlier concepts of coupling terms in order to propose a bonded-only treatment for 1-4 interactions, by implementing the coupling terms and developing the necessary protocols for generating and fitting QM data within Q-Force toolkit.25 Q-Force is a framework that enables a systematic, reproducible, and extensible approach to automated force field parameterization based on QM calculations which we harness for defining bonded coupling terms across 1-2, 1-3, and 1-4 coupling terms. We benchmark our results against traditional force fields, including OPLS and Sage, the polarizable force field AMOEBA, and the semi-empirical method xTB 26. Our findings reveal substantial improvements in the accuracy of forces and energy surfaces, demonstrating that a bonded-only treatment of 1-4 interactions can effectively overcome the limitations of hybrid approaches. Additionally, by decoupling the parameterization of torsional and non-bonded terms, our method simplifies force field development, providing a clear and streamlined protocol that can be readily adopted by next-generation of both two-body and advanced force fields. To demonstrate the promise of this approach, we replace the 1-4 nonbonded interactions of AMBER ff14SB, CHARMM36, and OPLS-AA with our bonded-only model for alanine dipeptide in the gas phase and under implicit solvent, and find excellent agreement with the QM data.
@@ 2 Theory and Methods
In this work, we derive a fully bonded approach for torsion-containing molecules and show that with the right torsion and torsion coupling terms, we can reproduce the PES without needing any 1-4 non-bonded interactions. We first consider bond and bond angle terms in their couplings derived in previous work.27 We then extend this to new torsion coupling terms by introducing them into Q-Force, which required a complete revamping of the QM data generation and the fitting procedures, which we also describe in this section.
@@ 2.1 Bond, angle, and bond-angle coupling terms
In earlier work, we demonstrated that with the correct selection of bond, angle, and coupling terms, we can reproduce the PES in small molecules with very high accuracy 27. We continue to exploit the functional forms for bond, angle, and bond-angle couplings for the molecules in Figure 1, which we briefly recapitulate here.
@@
In Q-Force, only the harmonic force constant $K_{b}$ is fitted, whereas $D_{e}$ is looked up from a table containing experimental bond dissociation energies 28. Coupling between bond lengths is captured by the bond-bond coupling term
@@
where $\theta_{0}$ is the equilibrium angle, and $K_{\theta}$ is the associated force constant.29 This cosine-based formulation introduces accurate anharmonicity to the potential without resorting to higher-order polynomials as is done in some force fields such as AMOEBA.10, 30 For angular coupling, the angle-angle interaction term is given by
@@ 2.2 Torsion terms and their coupling
For flexible torsions, which are torsions with multiple minima, we use an adapted version of the CFF93 model23, 24. Dihedral torsions are described by the Fourier functional form
@@
2.3 Parameterization and validation strategy
To validate our approach, we evaluate its performance across a diverse set of molecular systems encompassed in two datasets: the ”flexible set” featuring molecules with flexible torsions (Figure 1a) and the ”rigid set” comprising of rigid molecules with either ring systems or double bonds (Figure 1b). An important choice made here is that we only selected molecules with no discernible non-bonded interactions beyond 1-4, so that we can study the performance of the 1-4 bonded terms without having to choose a specific non-bonded FF, which could affect the interability of the results.
The 1-4 force fields were parameterized using the latest version of Q-Force25, which will be discussed in detail in a separate paper. However, below are the steps that are relevant for this work:
  • CREST 31 is used to generate all conformers of the molecule.
@@ 2.3 Parameterization and validation strategy
2.4 Protocol for 2D Torsion scans
@@
  • TorsionDrive32 using xTB GFN2 was used to generate optimized structures at each $\phi$ and $\psi$ point. All angles were held fixed from $165^{\circ}$ to $180^{\circ}$ in $15^{\circ}$ increments. (See Supplementary Figure S4)
@@
  • All 1-5 and beyond interactions from AMBER ff14SB, CHARMM36, and OPLS-AA were subtracted from the QM surface, and Q-Force used to fit the 1-2, 1-3, and 1-4 bonded terms.
  • The OpenMM software platform33 was used to perform energy minimizations for all force fields except for OPLS-AA for which GROMOS software package34 was used since OpenMM does not support OPLS-AA natively.
2.5 QM reference and other benchmark methods
QM reference data was generated using the $\omega$-B97X-V 35 density functional, one of the most accurate hybrid-GGA functionals available 36, 37, in conjunction with the def2-TZVPD 38, basis set. This combination was selected as a good compromise between accuracy and cost. The parameterization protocols in Q-Force, however, are agnostic to the choice of QM method, allowing for the flexibility to adopt more accurate or computationally efficient methods depending on the specific needs of the user or application. Q-Chem software was used for all QM calculations 39.
To benchmark Q-Force against other FFs, we parameterized three additional FFs using well-established tools. The OPLS force field, a widely used standard FF, was parameterized using LigParGen 40. Sage, a modern and QM-based FF developed within the Open Force Field Initiative, was parameterized using OpenForceField tools 41. For a comparison with an advanced FF, we included AMOEBA, a polarizable FF that incorporates various bonded coupling terms10, parameterized using the Poltype2 toolkit 42. These three FFs have 1-4 non-bonded interactions that are scaled by different constants. The xTB GFN2 semi-empirical model26 was also used as an additional benchmark to compare its performance against the FFs, as it uniquely occupies the space between FFs and QM methods in terms of both cost and accuracy. The OpenMM software platform was used for all FF evaluations.
@@ 2.5 QM reference and other benchmark methods
We recognize that comparing Q-Force FF, parameterized against the $\omega$-B97X-V reference, with other FFs optimized against different QM methods or objectives, does not constitute a perfectly fair comparison. For instance, Sage is parameterized against B3LYP-D3BJ with DZVP basis set. In order to estimate possible deviations arising from differences in QM references, we included the B3LYP-D3BJ 43, 44, 45 with TZVP basis set as a benchmark. Any error arising from differences in QM references should have similar magnitude to the deviation between these two QM methods. On the other hand, for AMOEBA, we adjusted Poltype settings so that any QM calculations it performs are done with $\omega$-B97X-V with the def2-TZVPD basis set.
@@ 3.1 Energy and force accuracy along the MD trajectory
The polarizable AMOEBA FF does better than the fixed charge FFs across energies and forces for the flexible set, but the results are more mixed for the rigid set. OPLS outperforms AMOEBA in energies while AMOEBA does slightly better with forces in this case. AMOEBA’s performance is particularly worse for ring systems, indicating that the automated parameterization strategy of Poltype2 toolkit 42 could be improved in this regard. The semi-empirical xTB’s performance for both energies and forces lies between the FFs and QM methods as expected, with $\sim$1 kcal/mol MAE in energies across both datasets.
@@ 3.3 Replacing 1-4 Non-bonded Interactions in Protein Force Fields
To evaluate the efficacy and validity of our new approach to treating 1-4 interactions for any type of force field, we consider the two dimensional $\phi/\psi$ scans of alanine dipeptide. Alanine dipeptide is considered to be a building block of protein models, and the ability to correctly calculate its Ramachandran surface is key to elucidating if a force field is well suited to model conformational changes in proteins. Furthermore, unlike the molecules from the flexible and rigid datasets, alanine dipeptide contains numerous 1-5 and beyond (1-5+) interactions. Hence we use Q-Force to refit the 1-2, 1-3, and 1-4 to the coupled bonded terms and embed them into the AMBER ff14SB, CHARMM36, and OPLS-AA 1-5+ interactions.
Figure 4A shows the QM relaxed $\phi/\psi$ surface of alanine dipeptide in the gas phase. This is compared to the Q-Force derived FFs with the bonded-only functional form, also in the gas phase, as seen in Figures 4B-D. The newly fitted maps reproduce the QM basin locations
@@
and barrier topography very well across the whole $(\phi,\psi)$ surface. Averaged across the 3 different force fields representing the 1-5+ interactions yields a mean absolute error (MAE) $\approx 0.3$ kcal/mole, root mean squared error (RMSE) $\approx 0.5$ kcal/mole, and absolute maximum error of $|\Delta E|\approx 2.25$ kcal/mol at the highest energy points on the map whose energy scale is $\sim$ 100 kcal/mol (see Supplementary Table S3). These deviations arise from the imperfect 1-5+ non-bonded potential from the fixed charge force fields that is used after the bonded fitting, and indicate that they are unable to capture all electronic effects from the QM reference such as polarization, charge transfer, and penetration due to missing many-body interactions.
As previous theory has shown46, the molecular origin of the $\alpha$-helical and left-handed helix minima in the dipeptide Ramanchandran plot is the stabilization that the reaction field from solvent imparts to the aligned peptide dipoles of these configurations, which is a repulsive interaction in the gas phase. This is born out as seen when comparing the gas phase QM result in Figure 4A with the QM and implicit solvent SMD model in Figure 4E. It also allows us to use Q-Force to create a bonded model but now in the presence of implicit solvent as seen in Figure 4F. In the solvated case the agreement with the QM-SMD surface is less good than the gas phase, with MAE $\approx 1.02$ kcal/mole, RMSE $\approx 1.33$ kcal/mole, and max—ΔE— $\approx 3.69$. The major contributor to the larger deviation is in the choice of solvent model employed, i.e., SMD at the QM level versus GBN2 at the MM level.

### v3

@@ 1 Introduction
The more common approach, employed by force fields such as OPLS 1988WLJorgensen, 2015MJRobertson, AMBER 1995WDCornell, 2004JWang, CHARMM 1998ADMacKerell, Vanommeslaeghe2010, the Open Force Field class of FFs Qiu2021, Boothroyd2023, and AMOEBARen2003, Ponder2010, is to use a single scaling factor to modulate 1-4 non-bonded interactions. While this method is computationally simple and easy to implement, it is suboptimal for a diverse set of molecular interactions for several reasons. A single scaling parameter cannot capture the nuanced physical phenomena governing 1-4 interactions across different chemical environments, leading to reduced accuracy in force field predictions. While GROMOS 1999WRPScott, 2011NSchmid incorporates unique 1-4 parameters in specific cases, it defaults to scaled interactions for most systems. While this tailored approach has the potential to improve accuracy for certain molecules or functional groups, it does not fully address the underlying limitations of the standard Lennard-Jones and Coulomb potentials. These functional forms inherently lack the correct physics to represent interactions at short distances characteristic of 1-4 pairs. Although torsional barriers can be improved by fine-tuning 1-4 parameters, this does not guarantee accurate forces or a well-behaved potential energy surface (PES), which are critical for reliable simulations.
@@
To address the inherent limitations of traditional methods for handling 1-4 interactions, we propose two possible pathways for improvement. The first involves treating 1-4 interactions entirely through non-bonded potentials with physically accurate functional forms. This would necessitate the development or adoption of advanced force fields where all non-bonded interactions, including those at short distances, are damped using physically motivated corrections, such as density overlap approaches kim1981dependence, wheatley1990overlap, van2016beyond. Force fields like HIPPO Rackers2021 and CMM Heindel2024, which incorporate density overlap to account for charge penetration and related effects, are promising candidates for this strategy. Such an approach could significantly reduce the reliance on torsional terms by inherently capturing the correct balance of interactions, including electrostatics, van der Waals forces, polarization, and charge transfer. However, while this approach holds promise, achieving the correct balance among these interaction components remains a considerable challenge. Subtle inaccuracies in modeling electrostatics, van der Waals interactions, or polarization effects could introduce errors that may still need to be compensated for by torsional terms. Additionally, the importance of polarization effects, particularly for 1-2 and 1-3 interactions, cannot be overlooked, as they can contribute significantly to the overall polarization energy at the 1-4 level Zgarbova2015, Tian2020. Therefore, even with an advanced non-bonded potential, some torsional corrections may still be necessary to address residual discrepancies and ensure an accurate representation of the energy landscape.
@@
The alternative approach involves treating the 1-4 interactions entirely as bonded terms. This bonded-only methodology not only circumvents the challenges associated with charge penetration effects discussed earlier but also decouples the non-bonded and torsional parameterization processes. This separation allows torsional terms to be directly optimized against quantum mechanical (QM) reference data without interference from non-bonded interactions. Such an approach necessitates handling the coupling between 1-4 interactions and 1-2 or 1-3 interactions—–effects that are traditionally, albeit imperfectly, addressed by non-bonded interactions in standard force fields. Fortunately, extensive research on coupled interaction terms, such as torsion-bond and torsion-angle coupling, provides a foundation for addressing these challenges. These terms were developed in the 1990s and implemented in force fields such as MM3 Allinger1989, Lii1989 and MMFF94 Halgren1996, Halgren1999, CFF93 Hwang1994, Maple1994. However, to our knowledge, all existing implementations still rely on non-bonded 1-4 interactions in addition to the coupling terms. Moreover, despite their potential, these coupling terms have so far not been widely adopted in common force fields. We believe the limited adoption of these coupling terms can be attributed to two primary challenges. First, there has been a lack of tools capable of systematically parameterizing this plethora of coupling terms. Second, generating the necessary QM reference data to calibrate these parameters accurately can also be a significant hurdle. Addressing these gaps requires an automated framework capable of deriving and optimizing these terms efficiently, as well as creating the necessary QM datasets.
@@
In this work, we revisit the earlier concepts of coupling terms in order to propose a bonded-only treatment for 1-4 interactions, by implementing the coupling terms and developing the necessary protocols for generating and fitting QM data within Q-Force toolkit.Sami2021 Q-Force is a framework that enables a systematic, reproducible, and extensible approach to automated force field parameterization based on QM calculations which we harness for defining bonded coupling terms across 1-2, 1-3, and 1-4 coupling terms. We benchmark our results against traditional force fields, including OPLS and Sage, the polarizable force field AMOEBA, and the semi-empirical method xTB Bannwarth2019. Our findings reveal substantial improvements in the accuracy of forces and energy surfaces, demonstrating that a bonded-only treatment of 1-4 interactions can effectively overcome the limitations of hybrid approaches. Additionally, by decoupling the parameterization of torsional and non-bonded terms, our method simplifies force field development, providing a clear and streamlined protocol that can be readily adopted by next-generation of both two-body and advanced force fields. To demonstrate the promise of this approach, we replace the 1-4 nonbonded interactions of AMBER ff14SB, CHARMM36, and OPLS-AA with our bonded-only model for alanine dipeptide in the gas phase and under implicit solvent, and find excellent agreement with the QM data.
@@ 2 Theory and Methods
In earlier work, we demonstrated that with the correct selection of bond, angle, and coupling terms, we can reproduce the PES in small molecules with very high accuracySami2024 which we briefly review here. In this work, we derive a fully bonded approach for torsion-containing molecules, with new torsion and torsion coupling terms introduced into Q-Force, which required a complete revamping of the QM data generation and the fitting procedures, which we also describe in this section.
@@ 2.1 Bond, angle, and bond-angle coupling terms
We continue to exploit the functional forms for bond, angle, and bond-angle couplings reported previouslySami2024 for the molecules in Figure 1, which we briefly recapitulate here.
@@
In Q-Force, only the harmonic force constant $K_{b}$ is fitted, whereas $D_{e}$ is looked up from a table containing experimental bond dissociation energies CRC_BDE_2016. Coupling between bond lengths is captured by the bond-bond coupling term
@@
where $\theta_{0}$ is the equilibrium angle, and $K_{\theta}$ is the associated force constant.Dateo1994 This cosine-based formulation introduces accurate anharmonicity to the potential without resorting to higher-order polynomials as is done in some force fields such as AMOEBA.Ponder2010, Liu2019 For angular coupling, the angle-angle interaction term is given by
@@ 2.2 Torsion terms and their coupling
For flexible torsions, which are torsions with multiple minima, we use an adapted version of the CFF93 modelHwang1994, Maple1994. Dihedral torsions are described by the Fourier functional form
@@
2.3 Q-Force parameterization strategy for small molecules
QM reference data was generated using the $\omega$-B97X-V Mardirossian2014 density functional, one of the most accurate hybrid-GGA functionals available Mardirossian2017, Goerigk2017, in conjunction with the def2-TZVPD Weigend2005, basis set. This combination was selected as a good compromise between accuracy and cost. Q-Chem software was used for all QM calculations Epifanovsky2021. The 1-4 force fields were parameterized with the QM data using the latest version of Q-ForceSami2021, which will be discussed in detail in a separate paper. However, below are the steps that are relevant for this work:
  • CREST Pracht2024 is used to generate all conformers of the small molecule.
@@ 2.3 Q-Force parameterization strategy for small molecules
To benchmark Q-Force against other FFs, we parameterized three additional FFs using well-established tools. The OPLS force field, a widely used standard FF, was parameterized using LigParGen Dodda2017. Sage, a modern and QM-based FF developed within the Open Force Field Initiative, was parameterized using OpenForceField tools Wagner2024. For a comparison with an advanced FF, we included AMOEBA, a polarizable FF that incorporates various bonded coupling termsPonder2010, parameterized using the Poltype2 toolkit Walker2022. These three FFs have 1-4 non-bonded interactions that are scaled by different constants. The xTB GFN2 semi-empirical modelBannwarth2019 was also used as an additional benchmark to compare its performance against the FFs, as it uniquely occupies the space between FFs and QM methods in terms of both cost and accuracy. The OpenMM software platform was used for all FF evaluations.
We recognize that comparing Q-Force FF for small molecules, parameterized against the $\omega$-B97X-V reference, with other FFs optimized against different QM methods or objectives, does not constitute a perfectly fair comparison. For instance, Sage is parameterized against B3LYP-D3BJ with DZVP basis set. In order to estimate possible deviations arising from differences in QM references, we included the B3LYP-D3BJ Becke1992, Lee1988, Grimme2011 with TZVP basis set as a benchmark. Any error arising from differences in QM references should have similar magnitude to the deviation between these two QM methods. On the other hand, for AMOEBA, we adjusted Poltype settings so that any QM calculations it performs are done with $\omega$-B97X-V with the def2-TZVPD basis set.
2.4 Protocol for 2D torsion scans for alanine dipeptide
@@
  • TorsionDriveQiu2020 using xTB GFN2 was used to generate optimized structures at each $\phi$ and $\psi$ point. All angles were held fixed from $165^{\circ}$ to $180^{\circ}$ in $15^{\circ}$ increments. (See Supplementary Figure S4)
@@
  • All 1-5 and beyond interactions from AMBER ff14SB, CHARMM36, and OPLS-AA were subtracted from the QM surface, and Q-Force used to fit the 1-2, 1-3, and 1-4 bonded terms described by Eqs (1)-(11). By 1-5 and beyond we refer to the Lennard Jones and Coulomb potentials used to treat atom pairs in the same molecule separated by 4 or more bonds.
  • The OpenMM software platformEastman2017 was used to perform energy minimizations for all force fields except for OPLS-AA for which GROMOS software packageChristen2005 was used since OpenMM does not support OPLS-AA natively.
@@ 3 Results
To test our approach for replacing 1-4 nonbonded interactions with torsions and their couplings, we evaluate its performance across a diverse set of small molecule systems encompassed in two datasets: the ”flexible set” featuring molecules with flexible torsions (Figure 1a) and the ”rigid set” comprising of rigid molecules with either ring systems or double bonds (Figure 1b). An important choice made here is that we only selected molecules with no discernible non-bonded interactions beyond 1-4, so that we can study the performance of the 1-4 bonded terms without having to choose a specific non-bonded FF, which could affect the interpretability of the results.
@@ 3.1 Energy and force accuracy along the MD trajectory
The polarizable AMOEBA FF does better than the fixed charge FFs across energies and forces for the flexible set, but the results are more mixed for the rigid set. OPLS outperforms AMOEBA in energies while AMOEBA does slightly better with forces in this case. AMOEBA’s performance is particularly worse for ring systems, indicating that the automated parameterization strategy of Poltype2 toolkit Walker2022 could be improved in this regard. The semi-empirical xTB’s performance for both energies and forces lies between the FFs and QM methods as expected, with $\sim$1 kcal/mol MAE in energies across both datasets.
@@ 3.3 Replacing 1-4 Non-bonded Interactions in Protein Force Fields
To evaluate the efficacy and validity of our new approach to treating 1-4 interactions for any type of force field, we consider the two dimensional $\phi/\psi$ scans of alanine dipeptide. Alanine dipeptide is considered to be a building block of protein models, and the ability to correctly calculate its Ramachandran surface is key to elucidating if a force field is well suited to model conformational changes in proteins. Furthermore, unlike the molecules from the flexible and rigid datasets, alanine dipeptide contains numerous 1-5 and beyond (1-5+) interactions.
We use Q-Force to fit the 1-2, 1-3, and 1-4 bonded terms and their couplings to the ab initio $\phi,\psi$ surface ($\omega$-B97X-V/def2-TZVPD) after subtracting the 1-5+ nonbonded interactions for each parent FF (OPLS, Amber, Charmm). It is important to note that we do not include the dihedral potentials from AMBER ff14sb, CHARMM36, and ff14sb, and the Coulomb and LJ parameters of the parent force field do not change. After fitting, the final
@@
Q-Force/FF model is the sum of bond, angle, and torsion terms and their couplings, ”embedded” into the unchanged non-bonded terms from each force field using their native parameters, but without any 1-4 nonbonded interactions.
Figure 4A shows the QM relaxed $\phi/\psi$ surface of alanine dipeptide in the gas phase. This is compared to the Q-Force derived FFs with the bonded-only functional form, also in the gas phase, as seen in Figures 4B-D. The newly fitted maps reproduce the QM basin locations and barrier topography very well across the whole $(\phi,\psi)$ surface without need for 1-4 non-bonded interactions. Averaged across the 3 different force fields representing the 1-5+ interactions, the newly fitted 1-2, 1-3, and 1-4 terms yields very good accuracy, with a mean absolute error (MAE) $\approx 0.3$ kcal/mole, root mean squared error (RMSE) $\approx 0.5$ kcal/mole, and absolute maximum error of $|\Delta E|\approx 2.25$ kcal/mol at the highest energy points on the map whose energy scale is $\sim$ 100 kcal/mol as seen in Supplementary Table S3. Supplementary Table S4 provides error metrics for the Ramachandran energy surfaces between Q-Force and the reference MM force field itself, and yields a MAE of $\approx 0.5$ kcal/mole, RMSE of $\approx 0.9$ kcal/mole, and $|\Delta E|\approx 4.0$ kcal/mol on the same large energy scale. These deviations arise from the imperfect 1-5+ non-bonded potential from the fixed charge FFs, and indicate that they are unable to capture all electronic effects from the QM reference such as polarization, charge transfer, and penetration due to missing many-body interactions.
As previous theory has shownShang1994, the molecular origin of the $\alpha$-helical and left-handed helix minima in the dipeptide Ramanchandran plot is the stabilization that the reaction field from water solvent imparts to the aligned peptide dipoles of these configurations, which is a repulsive interaction in the gas phase. In addition, the Simmerling group used implicit solvent models with QM to parameterize the $\phi,\psi$ plots of amino acids to deal with this polarization issue.Tian2020 This is born out as seen when comparing the gas phase QM result in Figure 4A with the QM and implicit solvent SMD model in Figure 4E. It also allows us to use Q-Force to create a bonded model but now in the presence of implicit solvent as seen in Figure 4F. In the solvated case the agreement with the QM-SMD surface is less good than the gas phase, with MAE $\approx 1.02$ kcal/mole, RMSE $\approx 1.33$ kcal/mole, and max—ΔE— $\approx 3.69$. The major contributor to the larger deviation is in the choice of solvent model employed, i.e., SMD at the QM level versus GBN2 at the MM level.